the lenght of a rectangle is five more than the width.If the perimeter of the rectangle is 18 meters,find the dimensions of the rectangle?
Answer by rfer(16322)

2w+2l=P
2w+2(w+5)=18
2w+2w+10=18
4w=8
w=2
w+5=7